### v2

# On the space complexity of online convolution

Abstract.
We study a discrete convolution streaming problem. An input arrives as a stream of numbers $z=(z_{0},z_{1},z_{2},\ldots)$, and at time $t$ our goal is to output $(Tz)_{t}$ where $T$ is a lower-triangular Toeplitz matrix. We focus on space complexity; the algorithm can store a buffer of $\beta(t)$ numbers in order to achieve this goal.
We characterize space complexity when algorithms perform continuous operations. The matrix $T$ corresponds to a generating function $G(x)$. If $G(x)$ is rational of degree $d$, then it is known that the space complexity is at most $O(d)$. We prove a corresponding lower bound; the space complexity is at least $\Omega(d)$. In addition, for irrational $G(x)$, we prove that the space complexity is infinite. We also provide finite-time guarantees. For example, for the generating function $\frac{1}{\sqrt{1-x}}$ that was studied in various previous works in the context of differentially private continual counting, we prove a sharp lower bound on the space complexity; at time $t$, it is at least $\Omega(t)$.
J.D.A. is supported by Providentia, a Data Science Distinguished Investigator grant from Novo Nordisk Fonden. A.Y. is supported by a DNRF chair grant. Both are partially supported by BARC
1. Introduction
Streaming algorithms process data as it arrives while maintaining only a limited buffer of past information. A fundamental problem in this context is the efficient computation of structured linear transformations. For a linear transformation $T$, given the stream $z=(z_{0},z_{1},\ldots)$, the goal at time $t$ is to compute $(Tz)_{t}$. We focus on the case that $T$ is a lower-triangular Toeplitz matrix. Such matrices correspond to discrete convolutions, they appear in a variety of applications, including signal processing, time series analysis, and differential privacy (see e.g. [6, 10, 14, 20] and references therein).
We suggest the following streaming model (a formal definition follows). The state $s^{(t)}$ of the algorithm at time $t$ is a buffer of $\beta(t)$ numbers. When the algorithm receives $z_{t+1}$, it updates the state $(s^{(t)},z_{t+1})\mapsto s^{(t+1)}$. The algorithm is correct at time $t$ if there is a way to recover $(Tz)_{t}$ from $s^{(t)}$. Our objective is to characterize the minimum space (buffer size) required to solve this task for a given $T$.
The complexity of this problem is naturally linked to the structure of the generating function $G$ associated with $T$. It is known [12] that when $G(x)$ is a rational function of degree $d$, the space complexity is at most $O(d)$. Prior to our work, no corresponding lower bound was known. We close the gap and establish a matching lower bound of $\Omega(d)$. Furthermore, we analyze the case of irrational $G(x)$ and prove that for this case, the space complexity is infinite. Our results classify the space complexity for infinite streams based on the rationality-degree of $G(x)$.
What about finite streams? For a finite stream of length $t$, we prove that certain generating functions, such as $G_{1/2}(x)=1/\sqrt{1-x}$, require space at least $\Omega(t)$. This resolves an open question in differentially private (DP) continual counting [14, 7], a central primitive in DP deep learning where space efficiency is critical [22, 11]. A leading-constant optimal algorithm for the problem [16, 19] relies on solving the streaming task for the corresponding $T=T[G_{1/2}]$. Unsuccessful attempts at solving $T[G_{1/2}]$ with small buffer size led to further research on space versus accuracy tradeoffs [3, 12, 4]. Our results vindicate this line of research, and carry insights for future work on streaming DP linear queries.
1.1. The model
For a sequence $a=(a_{0},a_{1},\ldots)\in\mathbb{R}^{\mathbb{N}}$ of real numbers, [^1] In this text =$\mathbb{N}=\{0,1,2,\ldots\}$; matrices and vectors are zero-indexed. denote by $T=T[a]$ the lower-triangular Toeplitz matrix
$T_{i,j}=\begin{cases}a_{i-j}&\text{if }i\geq j,\\ 0&\text{if }i<j.\end{cases}$
Given the stream $z=(z_{0},z_{1},z_{2},\ldots)\in\mathbb{R}^{\mathbb{N}}$, at time $t\in\mathbb{N}$ we want to compute
$(Tz)_{t}$
which depends only on the history $z_{0},\ldots,z_{t}$. For example, for $a=(1,1,1,\ldots)$, we get the problem of releasing all prefix-sums on the input, which is also known as continual counting.
This work focuses on the space complexity of this streaming problem. Because we are working over the real numbers, we suggest the following model.
Definition (streaming model).
A streaming algorithm is a sequence of maps $m^{(t)}$ and $u^{(t)}$ as follows. Let $\beta:\{-1\}\cup\mathbb{N}\to\mathbb{N}$ be a buffer size function with $\beta(-1)=1$. Denote the state at time $t$ by $s^{(t)}\in\mathbb{R}^{\beta(t)}$. Initially, the state is $s^{(-1)}=0\in\mathbb{R}$. Upon receiving $z_{t}$, the algorithm updates its internal state:
$s^{(t)}=u^{(t)}(s^{(t-1)},z_{t}),$
where $u^{(t)}:\mathbb{R}^{\beta(t-1)}\times\mathbb{R}\to\mathbb{R}^{\beta(t)}$. The algorithms then outputs
$m^{(t)}(s^{(t)})$
where $m^{(t)}:\mathbb{R}^{\beta(t)}\to\mathbb{R}$. We assume that the update rules are continuous; for each $t$, the map $u^{(t)}$ is a continuous map. The algorithm solves the streaming problem defined by $a$ if for every $z$ and every $t$,
$m^{(t)}(s^{(t)})=(Tz)_{t}$
where $T=T[a]$.
The model above assumes that the update rules are continuous (the output maps $m^{(t)}$ are not assumed to be continuous). For computations over $\mathbb{R}$, this assumption is natural and allows to use the algebraic operations +,×+$\times$+ , × as well as any other continuous function like trigonometric functions, max operations and exponential functions. Continuous computations, however, are restricted and can not perform “arbitrary operations”.
Remark.
A continuity restriction on the update rules is necessary because without it every problem can be solved with a buffer size of two. The reason is that we can encode $z=(z_{0},\ldots,z_{t})\in\mathbb{R}^{t+1}$ using a pair $(r,x)\in\mathbb{R}\times[0,1]$ as follows. Let $\varphi:[0,1]\to[0,1]^{t+1}$ be a continuous space-filling curve. For each non-zero $z\in\mathbb{R}^{t+1}$, let $r=\|z\|_{2}$ and let $x\in[0,1]$ be so that $\varphi(x)=\frac{z}{r}$ (for $z=0$, define $r=0$ and $x=0$). The map $z\mapsto(r,x)$ is not continuous, but we can reconstruct $z$ from $(r,x)$ using the continuous map $(r,x)\mapsto r\cdot\varphi(x)$.
We think of the buffer size $\beta$ as the memory cost of the algorithm. As is standard in computational complexity theory, algorithms have costs and problems have complexities.
Definition.
For a time bound $\tau$, define $\mathsf{space}_{a}(t;\tau)$ as the minimum value $b$ so that there is a streaming algorithm with buffer size $\beta$ that computes $a$ correctly up to time $\tau$ and $\max_{t^{\prime}\leq t}\beta(t^{\prime})\leq b$. For infinite streams, the space complexity $\mathsf{space}_{a}(t)$ is defined as $\mathsf{space}_{a}(t)=\mathsf{space}_{a}(t;\infty)$.
We start by observing the following basic property of algebraic algorithms (we do not rely on it, but we find it informative). An algorithm is called algebraic if for all $t$, the maps $m^{(t)}$ and $u^{(t)}$ are polynomial maps.
Lemma 1 (algebraic algorithms are linear).
If the algorithm defined by the maps $\{(u^{(t)},m^{(t)}):t\in\mathbb{N}\}$ always outputs a linear function of $z$, then there is an algorithm that uses only linear maps that computes the same function of $z$.
1.2. Generating functions
It is convenient and useful to think of the sequence $a$ as a generating function. The sequence $a$ corresponds to the (ordinary) generating function $G=G[a]$ defined by
$G(x)=\sum_{k=0}^{\infty}a_{k}x^{k}.$
This expression is just a formal infinite sum. For more background on generating functions, see e.g. [24, 28].
Let us consider some examples. The first example is the continual counting generating function
$G_{1}(x)=\sum_{k=0}^{\infty}x^{k}.$
The space complexity of $G_{1}$ is $1$ because the algorithm can just update the memory by adding the observed $z_{t}$. A second example is the “exponential counting” function
$G_{e}(x)=\sum_{k=0}^{\infty}2^{k}x^{k}.$
The space complexity of $G_{e}$ is $1$ as well (multiply the memory by two, and then add the observed $z_{t}$). The third example is
$G_{1/2}(x)=\sum_{k=0}^{\infty}2^{-2k}\binom{2k}{k}x^{k};$
motivation for studying this function comes from continual counting in the DP framework [19, 16]. The fourth example is
$G_{C}(x)=\sum_{k=0}^{\infty}C_{k}x^{k},$
where $C_{k}$ is the $k$’th Catalan number. The trivial upper bound for the space complexities of both $G_{1/2}$ and $G_{C}$ is $\mathsf{space}(t)\leq t+1$. Is there a clever way to significantly reduce the buffer size? It turns out that the property that qualitatively governs the space complexity is rationality.
Definition.
A polynomial $Q(x)$ of degree $d$ is an expression of the form $Q(x)=\sum_{k=0}^{d}q_{k}x^{k}$ where $q_{d}\neq 0$. The polynomial $Q(x)$ is invertible if $q_{0}\neq 0$. The generating function $G$ is a rational function of degree at most $d$ if there are polynomials $P,Q$ with $Q$ invertible so that $Q(x)\cdot G(x)=P(x)$ where $\deg(Q)\leq d$ and $\deg(P)\leq d-1$. We say that $G$ has degree $d$ if it is of degree at most $d$ but not of degree at most $d-1$. We say that $G$ is irrational if it is not rational (of any degree).
Let us go back to the four examples we introduced:
$G_{1}(x)=\frac{1}{1-x},$
$G_{e}(x)=\frac{1}{1-2x},$
$G_{1/2}(x)=\frac{1}{\sqrt{1-x}},$
$G_{C}(x)=\frac{1-\sqrt{1-4x}}{2x}.$
We see that $G_{1},G_{e}$ are degree one rational functions; as simple as it gets. What about $G_{C}$ and $G_{1/2}$?
Claim.
Both $G_{1/2}$ and $G_{C}$ are irrational.
1.3. Results
Our main results are that the degree of rationality of $G$ determines the space complexity of the corresponding streaming problem.
Theorem 2.
If $G[a]$ is a rational function of degree $d$, then for $t\geq d$,
$\mathsf{space}_{a}(t)=d$
and for $t<d$,
$\mathsf{space}_{a}(t;t+d-1)=t+1.$
Theorem 3.
If $G[a]$ is irrational, then for all $t\in\mathbb{N}$,
$\mathsf{space}_{a}(t)=t+1.$
Theorem 3 holds just for infinite streams. Irrationality is insufficient to derive lower bounds for finite streams because for every $\tau$, there exists [^2] Consider for example $G(x)=G_{C}(x)+\sum_{k=0}^{\tau}(1-C_{k})x^{k}$. an irrational $G[a]$ so that $\mathsf{space}_{a}(t;\tau)=1$. The theorem does not provide any information on space complexity for a finite time horizon ($\tau<\infty$).
The analysis of finite time horizons is done by studying the corresponding Hankel matrix (see Section 3). Hankel matrices were studied in combinatorics and in the analytic theory of continued fractions [27]. The properties of Hankel matrices allow to prove that $G_{1/2}$ is essentially as hard as it gets even for finite times.
Theorem 4.
For all $t\in\mathbb{N}$,
$\mathsf{space}_{G_{1/2}}(t;2t)=t+1.$
The theorems reveal “Diophantine approximation like” properties that are related to the space complexity of streaming algorithms. Diophantine approximation studies “how well can a real number $\alpha$ be approximated by rationals?” If $\alpha$ is rational, then it has a perfect representation. If $\alpha$ is irrational, it can only be approximated. There are standard ways to measure the cost of such approximations. The “hardest to approximate” real numbers turn out to be irrational algebraic numbers (e.g. the Thue–Siegel–Roth theorem). For example, the golden ratio $\frac{1+\sqrt{5}}{2}$ has “high complexity” in this model (notice the similarity to $G_{C})$. In our context, space complexity corresponds to “distance from rationality”. Rational functions have constant complexity, and irrational ones have infinite complexity. In addition, like the golden ratio, the algebraic irrationals $G_{1/2}$ and $G_{C}$ are “as far as possible from the rationals”.
1.4. Applications
The above results also carry algorithmic implications. Continual counting is about computing prefix-sums of the stream. In our setting, it corresponds to $T_{1}=T[(1,1,\ldots)]$. Differentially private (DP) continual counting [14, 7] solves the same problem, but also requires the streaming algorithm $\mathcal{A}$ to satisfy differential privacy [13, 15].
Throughout this section, we only consider the approximate DP version of the problem, which sees the most practical use. We also only consider the case where an upper bound on the length of the stream is known. [^3] A technique in [7] allows for removing the assumption of a bounded stream at the cost of worse guarantees, but complicates the exposition.
Most DP algorithms for continual counting are based on adding correlated noise to the exact output and have the form [11]:
$\mathcal{A}(z)=T_{1}z+Ly,$
where $L$ is a lower-triangular matrix and $y$ is a random Gaussian vector sampled from $\mathcal{N}(0,\sigma^{2}I)$. The noise parameter $\sigma$ is determined by the privacy parameters and by the matrix $L$. Observe that $\mathcal{A}(z)$ is essentially running two instances of our streaming problem: upon receiving $z_{t}$, it generates a sample $y_{t}$ and outputs $(T_{1}z)_{t}+(Ly)_{t}$. The space complexity of $T_{1}$ is $1$, and so the space complexity of $\mathcal{A}(z)$ is determined by $L$. To provide the DP guarantee, the computation of $(Ly)_{t}$ must be exact.
Space complexity matters for this problem. Part of the recent interest in DP continual counting is motivated by private machine learning [22, 11, 9, 8, 25]. The main observation is that unconstrained stochastic gradient descent updates with constant learning rate can be viewed as computing prefix sums on adaptively chosen gradient vectors. In other words, the training of a model reduces to running one instance of continual counting per model dimension. Because the model dimension often dominates the number of weight updates [12], a linear space complexity is prohibitive in practice.
The choice of the matrix $L$ has a crucial impact on the DP properties of $\mathcal{A}$. Choosing $L=\sqrt{T_{1}}=T_{1/2}$ leads to near-optimal guarantees [16, 19]. [^4] Very recently, an explicit construction of a non-Toeplitz $L$ improving lower order terms was given in [18]. No space-efficient implementation for this $L$ is known. However, attempts to implement $T_{1/2}$ in sublinear space were unsuccessful. This drove further research into other matrices $L$ and into accuracy versus space tradeoffs [3, 12, 4, 17]. The best known bounds for this trade-off [12] is achieved for $L=T[\tilde{G}_{1/2}]\approx T_{1/2}$, where $\tilde{G}_{1/2}$ is a rational-function approximation of $G_{1/2}$ implemented with memory cost equal to the degree of $\tilde{G}_{1/2}$. Theorem 2 implies that this implementation of $T[\tilde{G}_{1/2}]$ is tight, and Theorem 4 motivates why matrices other than $T_{1/2}$ had to be considered for the problem. In a nutshell, there are no sublinear space algorithms for $T_{1/2}$. Any implementation of $\mathcal{A}$ with $L=T_{1/2}$ for $n$ steps, restricted to continuous operations, requires space $\Omega(n)$.
Additional motivation from private machine learning led to studying the problem of DP continual counting with momentum and weight decay [23]. Translating this problem to our setting, there is a momentum parameter $10\leq\mu<$ 1 and a weight decay parameter and $0<\lambda\leq 1$. To avoid degenerate cases, we can assume that $0\leq\mu<\lambda\leq 1$. The streaming problem now corresponds to the generating function
$G_{\lambda,\mu}=\frac{1}{(1-\lambda x)(1-\mu x)}.$
In [23], a DP streaming algorithm of the form
$\mathcal{A}_{\lambda,\mu}(z)=T_{\lambda,\mu}z+Ly$
was presented, achieving asymptotically optimal error for $L=T[\sqrt{G_{\lambda,\mu}}]$. Our methods yield optimal lower bounds for the space complexity of this task as well.
Theorem 5.
For $0\leq\mu<\lambda\leq 1$ and for all $t\in\mathbb{N}$,
$\mathsf{space}_{\sqrt{G_{\lambda,\mu}}}(t;2t)\geq t-4.$
There is additional literature on DP continual counting with arbitrary weights [5, 20, 18]. Most algorithms in this domain are also based on adding correlated noise $Ly$, and whenever $L$ is Toeplitz our techniques and results can be leveraged.
Remark.
In a couple of places below we reuse notation. For example, $\mu$ denotes the momentum parameter above and in one of the proofs below it denotes a monomial. The meaning is clear from the context.
2. Algebraic algorithms are linear
An algebraic streaming algorithm is defined by a sequence of polynomial maps $u^{(t)}$ and $m^{(t)}$. Our first observation is that we can assume (without loss of generality) that the maps $u^{(t)}$ and $m^{(t)}$ are linear. For a multivariate polynomial $p$ and $d\in\mathbb{N}$, denote by $\mathop{\mathcal{H}_{d}}(p)$ the homogeneous part of degree $d$ in $p$ (it is a linear functional).
Lemma 6.
Let $f:\mathbb{R}^{n}\to\mathbb{R}^{k}$ and $g:\mathbb{R}^{k}\to\mathbb{R}$ be two polynomial maps. Write $f=(f_{1},\ldots,f_{k})$, where each $f_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is a polynomial map. Then, there are linear maps $L_{f}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ and $L_{g}:\mathbb{R}^{k}\to\mathbb{R}$ so that
$L_{g}\circ L_{f}=\mathop{\mathcal{H}_{1}}(g\circ f).$
In addition, $(L_{f})_{i}=\mathop{\mathcal{H}_{1}}(f_{i})$ for all $i\in[k]$.
Proof.
For polynomials over $\mathbb{R}^{n}$, equality as functions is equivalent to equality as formal polynomials. Define the affine map $A_{f}:\mathbb{R}^{n}\to\mathbb{R}^{k}$ by
$(A_{f})_{i}=\mathop{\mathcal{H}_{0}}(f_{i})+\mathop{\mathcal{H}_{1}}(f_{i}).$
Every monomial in $f_{i}(x)-(A_{f})_{i}(x)$ has degree at least two. So, for every $k$-variate monomial $\mu$, it holds that
$\mathop{\mathcal{H}_{1}}(\mu(f(x)))=\mathop{\mathcal{H}_{1}}(\mu(f(x)-A_{f}(x)+A_{f}(x)))=\mathop{\mathcal{H}_{1}}(\mu(A_{f}(x))).$
It follows that
$\mathop{\mathcal{H}_{1}}(g(f(x)))=\mathop{\mathcal{H}_{1}}(g(A_{f}(x))).$
For every $k$-variate monomial $\mu$, there is a polynomial $q=q_{f,\mu}$ so that
$\mathop{\mathcal{H}_{1}}(\mu(A_{f}(x)))=\mathop{\mathcal{H}_{1}}(q(L_{f}(x)))$
where $(L_{f})_{i}:=\mathop{\mathcal{H}_{1}}(f_{i})$. By linearity, there is a polynomial $r=r_{f,g}$ so that
$\mathop{\mathcal{H}_{1}}(g(A_{f}(x)))=\mathop{\mathcal{H}_{1}}(r(L_{f}(x))).$
If $\mu$ is a monomial of (total) degree larger than one, then $\mu(L_{f}(x))$ has degree larger than one (unless it is zero). It follows that
$\mathop{\mathcal{H}_{1}}(r(L_{f}(x)))=L_{g}(L_{f}(x))$
where $L_{g}=\mathop{\mathcal{H}_{1}}(r)$. ∎
Proof of Lemma 1.
The update of the state at time $t$ is
$s^{(t)}=u^{(t)}(s^{(t-1)},z_{t}).$
Let $L_{u}^{(t)}$ be the linear map obtained with $g=u^{(t)}$ in Lemma 6. Inductively define states $\sigma^{(t)}$ by $\sigma^{(-1)}=0$ and
$\sigma^{(t)}=L_{u}^{(t)}(\sigma^{(t-1)},z_{t}).$
It follows by induction and Lemma 6 that
(1) $\mathop{\mathcal{H}_{1}}(s^{(t)})=L_{u}^{(t)}(\mathop{\mathcal{H}_{1}}(s^{(t-1)}),z_{t})=L_{u}^{(t)}(\sigma^{(t-1)},z_{t})=\sigma^{(t)}.$
By assumption, the output at time $t$ is a linear function:
$m^{(t)}(s^{(t)})=\mathop{\mathcal{H}_{1}}(m^{(t)}(s^{(t)})).$
By Lemma 6 with $g=m^{(t)}$ and using (1), there is a linear map $L_{m}^{(t)}$ so that
$\mathop{\mathcal{H}_{1}}(m^{(t)}(s^{(t)}))=L_{m}^{(t)}(\sigma^{(t)}).\qed$
3. A lower bound mechanism
In this section, we develop a mechanism for proving lower bounds on the space complexity of a sequence $a$ or a Toeplitz matrix $T=T[a]$. For the lower bound, we shall also employ the Hankel matrix $H=H[a]$ defined by
$H_{i,j}=a_{i+j}.$
For $I,J\geq 0$, we denote by $H^{(I,J)}$ the $I+1\times J+1$ matrix that corresponds to positions $0\leq i\leq I$ and $0\leq j\leq J$ in $H$.
Lemma 7 (small buffer, small rank).
For all $t,I\in\mathbb{N}$,
$\mathsf{space}_{a}(t;t+I)\geq\mathrm{rank}(H^{(I,t)})$
where $H=H[a]$.
The lemma relates space complexity, which is a “complicated notion”, to the rank of a matrix, which is a notion that we understand pretty well.
Remark.
A variant of the lemma holds in a more general framework than the one we consider. It holds for general lower-triangular matrices and not just for lower-triangular Toeplitz matrices. For every matrix $A$, the buffer size can be bounded from below by the rank of some submatrix of $A$.
The lower bound is based on a basic topological property; the Borsuk-Ulam theorem. Denote by $\mathbb{S}^{r-1}$ the unit sphere in $\mathbb{R}^{r}$:
$\mathbb{S}^{r-1}=\{x\in\mathbb{R}^{r}:\|x\|_{2}=1\}.$
Theorem (Borsuk-Ulam).
If $f:\mathbb{S}^{r-1}\to\mathbb{R}^{r-1}$ is continuous then there is $x\in\mathbb{S}^{r-1}$ so that $f(x)=f(-x)$.
Proof of Lemma 7.
Consider the restricted setting where the incoming stream $z$ is so that $z_{t^{\prime}}=0$ for all $t^{\prime}>t$, and we are only interested in the outputs $(Tz)_{t^{\prime}}$ for $t\leq t^{\prime}\leq t+I$. Accordingly, for the rest of the proof, assume that $z$ is zero after position $t$, or equivalently that $z\in\mathbb{R}^{t+1}$. Denote by $T^{\prime}$ the first $t+1$ columns of $T$ so that $Tz=T^{\prime}z$.
Denote by $f(z)\in\mathbb{R}^{\beta(t)}$ the state of the memory at time $t$ when the input is $z\in\mathbb{R}^{t+1}$. The map $f$ is continuous by assumption on the algorithm. Denote by $M$ the submatrix of $T^{\prime}$ formed from row indices $t,t+1,\ldots,t+I$; it defines a linear map from $\mathbb{R}^{t+1}$ to $\mathbb{R}^{I+1}$. By definition of rank, there is a vector space $V\subset\mathbb{R}^{t+1}$ of dimension $r=\mathrm{rank}(M)$ so that $Mv\neq 0$ for every non-zero $v\in V$. The space $V$ is isomorphic to $\mathbb{R}^{r}$. Let $S$ be a copy of the sphere $\mathbb{S}^{r-1}$ in $V$. For every $v\in S$ we have $Mv\neq 0$.
Denote by $g(z)\in\mathbb{R}^{I+1}$ the output of the algorithm at times $t,t+1,\ldots,t+I$ on input $z$. By the structure of the algorithm, $g$ can be factored as $g=h\circ f$ where $h:\mathbb{R}^{\beta(t)}\to\mathbb{R}^{I+1}$ (we do not know that $h$ is continuous). Assume towards a contradiction that $\beta(t)<r$. By the Borsuk-Ulam theorem, there is a $v\in S$ so that $f(v)=f(-v)$ and consequently $g(v)=g(-v)$. By the correctness criterion,
$Mv=g(v)=g(-v)=-Mv$
so $Mv=0$ which is a contradiction.
Because $H^{(I,t)}$ is a column permutation of $M$, the proof is complete. ∎
4. Asymptotic complexity
In this section, we show that the space complexity of infinite streams is characterized by the rational function degree of the generating function.
Compactness
Lemma 8.
Let $M$ be a $d\times\mathbb{N}$ real matrix with row rank $d$. For each $j\in\mathbb{N}$, let $M[j]$ be the $d\times j+1$ matrix of the first $j+1$ columns of $M$. Then, there is some $j\in\mathbb{N}$ so that $M[j]$ has rank $d$.
Proof.
Assume towards a contradiction that for all $j$, the matrix $M[j]$ is singular. That is, there is a $v[j]$ in the unit sphere of $\mathbb{R}^{d}$ so that $v[j]M[j]=0$; in words, it is in the left kernel of $M[j]$. The unit sphere is compact so by going to a subsequence we can assume that $v[j]$ tends to a limit $v_{*}\neq 0$.
Fix $j_{0}\in\mathbb{N}$, and denote by $m\in\mathbb{R}^{d}$ the $j_{0}$’th column in $M$. For all $j>j_{0}$, we have $\langle v[j],m\rangle=0$. As $j\to\infty$, we have $\langle v_{*},m\rangle=\langle v_{*}-v[j],m\rangle\to 0$. So, $\langle v_{*},m\rangle=0$.
It follows that $v_{*}$ is in the left kernel of $M$, which is a contradiction. ∎
Hankel matrices
We recall the following theorem proved by Kronecker; see e.g. [26, 2] and references therein.
Theorem 9.
Let $a\in\mathbb{R}^{\mathbb{N}}$. The generating function $G=G[a]$ is rational of degree $d$ iff $H=H[a]$ has rank $d$ and $H^{(d-1,d-1)}$ has rank $d$. The function $G$ is irrational iff the rank of $H$ is infinite.
Claim 10.
Let $a\in\mathbb{R}^{\mathbb{N}}$ and let $H=H[a]$. If the rank of $H$ is $d$, then for every $t\leq d$, the first $t$ columns in $H$ are linearly independent.
Proof.
If the first $t$ columns of $H$ are linearly dependent, then some column $i\leq t$ is a linear combination of columns $<i$. Because $H$ is an infinite Hankel matrix, all columns after column $i$ are also spanned by columns $<i$ (with the same coefficients). ∎
Rational functions
Proof of Theorem 2.
Assume that $G=G[a]$ is a rational function of degree $d$ and let $H=H[a]$ be its Hankel matrix. It is known that because $G[a]$ is rational of degree $d$, it holds that $\mathsf{space}_{a}(t)\leq d$ for all $t$; see (1.8) and Lemma 3.2 in [12]. Their algorithm as stated requires $\beta(0)=d$, but it is possible to implement it with $\beta(t)=\min\{t+1,d\}$. For $t<d$, let $s^{(t)}$ be a copy of the input stream $s^{(t)}=(z_{0},\ldots,z_{t})$ with the obvious output map $m^{(t)}$. For $t\geq d$, let $s^{(d)}$ be the corresponding state from [12] and use their (fixed) maps $\{(u^{(t)},m^{(t)}):t\geq d\}$ for all future computations. It follows that $\mathsf{space}_{a}(t)\leq\min\{t+1,d\}$.
Now, let $t<d$. By Theorem 9, the matrix $H^{(d-1,d-1)}$ has full rank. In particular, $\mathrm{rank}(H^{(d-1,t)})=t+1$. Invoking Lemma 7, we get $\mathsf{space}_{a}(t;t+d-1)\geq t+1$.
Irrational functions
Proof of Theorem 3.
Because $G[a]$ is irrational, the Hankel matrix $H[a]$ has infinite rank. By 10, for all $t\in\mathbb{N}$, the first $t+1$ columns in $H$ are linearly independent. Fix $t$ and let $M$ be the first $t+1$ columns of $H$. By compactness, there is some $j$ so that $H^{(j,t)}$ has rank $t+1$. By Lemma 7, we know that $\mathsf{space}_{a}(t)\geq t+1$. The upper bound $\mathsf{space}_{a}(t)\leq t+1$ is trivial. ∎
5. Finite time horizons
Asymptotic space complexity can be understood via degrees as rational functions. The finite time horizon case for irrational functions can be understood via deeper properties of the corresponding Hankel matrices.
Preliminaries
In this section, we explain some methods for comparing the ranks of different Hankel matrices. We shall use the following simple properties. If $f(x),g(x)$ are generating functions and $T[f],T[g]$ are the corresponding Toeplitz matrices then
$T[f\cdot g]=T[f]\cdot T[g]$
and
$T[f+g]=T[f]+T[g].$
Lemma 11.
Let $f,g$ be generating functions and let $\alpha,\beta,\gamma\in\mathbb{R}$ be so that $f_{0}\neq\alpha$, so that $g_{0}\neq 0$ and so that
$fg=\alpha g+\beta xg+\gamma.$
Then, for every $d\in\mathbb{N}$, the rank of $(H[f])^{(d,d)}$ is at least the rank of $(H[g])^{(d,d)}$ minus three. In addition, when $\beta=0$, the rank of $(H[f])^{(d,d)}$ is at least the rank of $(H[g])^{(d,d)}$ minus two.
Proof.
Fix $d$, and denote by $T^{\prime}[h]$ the principle $2d\times 2d$ submatrix of $T[h]$. Because the “upper half” of $T[h]$ is zero,
$T^{\prime}[f\cdot g]=T^{\prime}[f]\cdot T^{\prime}[g].$
Therefore, by assumption,
$T^{\prime}[f-\alpha-\beta x]\cdot T^{\prime}[g]=T^{\prime}[\gamma].$
The matrix $T^{\prime}[h]$ can be partitioned into four equal block of size $d\times d$ as
$T^{\prime}[h]=\left[\begin{array}[]{cc}T_{11}[h]&T_{12}[h]\\ T_{21}[h]&T_{22}[h]\\ \end{array}\right].$
By looking at the lower-left block, $T_{21}$, we see that
$T_{21}[f-\alpha-\beta x]\cdot T_{11}[g]+T_{22}[f-\alpha-\beta x]\cdot T_{21}[g]=T_{21}[\gamma]=0,$
because $T_{21}[\gamma]$ is zero. The matrices $T_{11}[g]$ and $T_{22}[f-\alpha-\beta x]$ have full rank because $f_{0}\neq\alpha$ and $g_{0}\neq 0$. The matrix
$M:=-T_{22}[f-\alpha-\beta x]\cdot T_{21}[g]\cdot(T_{11}[g])^{-1}$
has the same rank as $T_{21}[g]$ and
$T_{21}[f-\alpha-\beta x]=M.$
So,
$T_{21}[f]=M+T_{21}[\alpha+\beta x]=M+T_{21}[\beta x].$
The matrix $T_{21}[\beta x]$ has at most one non-zero entry (so its rank is at most one). The matrix $T_{21}[h]$ is obtained from $(H[h])^{(d,d)}$ by first deleting the first row and first column, and then reversing the rows. The rank of $(H[f])^{(d,d)}$ is hence at least the rank of $T_{21}[f]$. The rank of $M$, which is equal to the rank of $T_{21}[g]$, is at least the rank of $(H[g])^{(d,d)}$ minus two. ∎
Examples
Let us start with the Catalan generating function $G_{C}$. Denote by $H_{C}$ the Hankel matrix that corresponds to the Catalan sequence. We note the following folklore lemma (see e.g. [1]).
Lemma 12.
For every $d\in\mathbb{N}$,
$\det(H_{C}^{(d,d)})=1.$
In particular, $H_{C}^{(d,d)}$ has full rank.
Denote by $H_{1/2}$ the Hankel matrix that corresponds to $G_{1/2}$. An $n\times n$ matrix has co-rank $r$ if its rank is $n-r$. The full rank of $(H_{C})^{(d,d)}$ can be used with Lemma 11 to show that $(H_{1/2})^{(d,d)}$ has a co-rank of at most four. The rank of $(H_{1/2})^{(d,d)}$ is, in fact, known to be full.
Lemma 13.
For every $d\in\mathbb{N}$, $(H_{1/2})^{(d,d)}$ has full rank.
Proof.
Proposition 8 in [1] states that $\det(H[1/\sqrt{1-4x}]^{(d,d)})=2^{d}$ implying the full rank of $(H[1/\sqrt{1-4x}])^{(d,d)}$. It follows that $(H[1/\sqrt{1-x}])^{(d,d)}$ also has full rank. ∎
Proof of Theorem 4.
The lower bound follows from Lemma 13 and Lemma 7. The upper bound $\mathsf{space}_{a}(t;2t)\leq t+1$ is trivial. ∎
Finally, recall the generating function
$G_{\lambda,\mu}=\frac{1}{(1-\lambda x)(1-\mu x)}$
for $0\leq\mu<\lambda\leq 1$. Our main lemma here is the following.
Lemma 14.
Let $b\neq c$ be non-zero complex numbers and let
$G(x)=\frac{8}{(b-c)^{2}}\frac{1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)}}{x^{2}}.$
For all $d\in\mathbb{N}$, the matrix $(H[G])^{(d,d)}$ has full rank.
Our argument relies on the following theorem by Junod (Theorem 2 in [21]).
Theorem 15.
Consider a generating function
$G(x)=\frac{1}{1-W(x)}$
with $W(0)=0$. Suppose that $w(x)=\frac{W(x)}{x}-W^{\prime}(0)$ satisfies
$w(x)=x(\alpha+\beta w(x)+\gamma w(x)^{2})$
for some parameters $\alpha\neq 0,\beta$ and $\gamma$ in $\mathbb{C}$. Then for any $d\in\mathbb{N}$:
$\det((H[G])^{(d,d)})=\alpha^{n(n+1)/2}\gamma^{n(n-1)/2}\enspace.$
Junod also remarks that $w(x)$ can be written as
$w(x)=\frac{F(x)-1}{xF(x)}-F^{\prime}(0)$
and as
$w(x)=\frac{1-\beta x-\sqrt{(1-\beta x)^{2}-4\gamma\alpha x^{2}}}{2\gamma x},$
when $\gamma\neq 0$.
Proof of Lemma 14.
To invoke Theorem 15, we will first prove that $W(x)=1-1/G(x)$ satisfies $W(0)=0$. We proceed:
$W(x)$ $=1-\frac{(b-c)^{2}}{8}\frac{x^{2}}{1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)}}.$
For some generating functions $R_{1},R_{2},R_{3}$, we can write
$\sqrt{(1-bx)(1-cx)}$
$=\sqrt{1-(b+c)x+bcx^{2}}$
$=1-\frac{(b+c)x}{2}+\frac{bcx^{2}}{2}-\frac{x^{2}(b+c-bcx)^{2}}{8}-\frac{x^{3}(b+c-bcx)^{3}}{16}+x^{4}R_{1}$
$=1-\frac{(b+c)x}{2}+x^{2}\Big{(}\frac{bc}{2}-\frac{(b+c-bcx)^{2}}{8}-\frac{x(b+c-bcx)^{3}}{16}+x^{2}R_{1}\Big{)}$
$=1-\frac{(b+c)x}{2}+x^{2}\Big{(}\frac{bc}{2}-\frac{b^{2}+c^{2}+2bc}{8}+\frac{2bc(b+c)}{8}x-\frac{(b+c)^{3}}{16}x+x^{2}R_{2}\Big{)}$
$=1-\frac{(b+c)x}{2}-\frac{x^{2}(b-c)^{2}}{8}\Big{(}1+\frac{b+c}{2}x+x^{2}R_{3}\Big{)}.$
So,
$W(x)$ $=1-\frac{1}{1+\frac{b+c}{2}x+x^{2}R_{3}},$
and we can conclude that $W(0)=0$. It also follows that $G^{\prime}(0)=\frac{b+c}{2}$ and by the remark above we can conclude
$w(x)$ $=\frac{G(x)-1}{xG(x)}-G^{\prime}(0)$
$=\frac{\frac{8}{(b-c)^{2}}\frac{1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)}}{x^{2}}-1}{\frac{8}{(b-c)^{2}}\frac{1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)}}{x}}-\frac{b+c}{2}$
$=\frac{8(1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)})-(b-c)^{2}x^{2}}{8x(1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)})}-\frac{b+c}{2}.$
We claim that
$\frac{8(1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)})-(b-c)^{2}x^{2}}{8x(1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)})}$
$\qquad=\frac{x-\frac{x}{2}(1-\frac{b+c}{2}x+\sqrt{(1-bx)(1-cx)})}{x^{2}}.$
Indeed, this is equivalent to
$x^{2}\Big{(}8\Big{(}1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)}\Big{)}-(b-c)^{2}x^{2}\Big{)}$
$\qquad=\Big{(}8x(1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)})\Big{)}$
$\qquad\qquad\cdot\Big{(}x-\frac{x}{2}(1-\frac{b+c}{2}x+\sqrt{(1-bx)(1-cx)})\Big{)}$
which is correct because
$\Big{(}8x(1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)})\Big{)}\frac{x}{2}\Big{(}1-\frac{b+c}{2}x+\sqrt{(1-bx)(1-cx)}\Big{)}$
$=4x^{2}\Big{(}1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)}\Big{)}\Big{(}1-\frac{b+c}{2}x+\sqrt{(1-bx)(1-cx)}\Big{)}$
$=4x^{2}\Big{(}\Big{(}1-\frac{b+c}{2}x\Big{)}^{2}-(1-bx)(1-cx)\Big{)}$
$=4x^{2}\Big{(}\Big{(}\frac{b+c}{2}x\Big{)}^{2}-bcx^{2})\Big{)}$
$=4x^{2}\Big{(}\frac{b^{2}+c^{2}+2bc}{4}x^{2}-bcx^{2})\Big{)}$
$=x^{2}(b-c)^{2}x^{2}.$
So,
$w(x)$ $=\frac{x-\frac{x}{2}(1-\frac{b+c}{2}x+\sqrt{(1-bx)(1-cx)})}{x^{2}}-\frac{b+c}{2}$
$=\frac{1-\frac{1}{2}(1-\frac{b+c}{2}x+\sqrt{(1-bx)(1-cx)})-\frac{(b+c)}{2}x}{x}$
$=\frac{1-\frac{b+c}{2}x-\sqrt{(1-bx)(1-cx)})}{2x}.$
As in the remark above,
$w(x)=\frac{1-\beta x-\sqrt{(1-\beta x)^{2}-4\gamma\alpha x^{2}}}{2\gamma x},$
with $\gamma=1$, with $\beta=\frac{b+c}{2}$ and with $\alpha=\frac{(b-c)^{2}}{16}$. It follows that the determinant of $(H[F])^{(d,d)}$ is non-zero. ∎
Denote by $H_{\lambda,\mu}$ the Hankel matrix that corresponds to $G_{\lambda,\mu}$.
Corollary 16.
For all $0\leq\mu<\lambda\leq 1$ and for all $d\in\mathbb{N}$, the matrix $(H_{\lambda,\mu})^{(d,d)}$ has co-rank at most five.
Proof.
Start with $f(x)=\sqrt{(1-\lambda x)(1-\mu x)}$ and with
$G(x)=\frac{1-\frac{\lambda+\mu}{2}x-f}{x^{2}}.$
Because
$f\cdot G$ $=\frac{(1-\frac{\lambda+\mu}{2}x)f-(1-\lambda x)(1-\mu x)}{x^{2}}$
$=-G-\lambda\mu+\frac{(\lambda+\mu)(1-f)}{2x}$
$=-G-\lambda\mu+\frac{(\lambda+\mu)(1-\frac{\lambda+\mu}{2}x-f)}{2x}+\frac{(\lambda+\mu)\frac{\lambda+\mu}{2}x}{2x}$
$=\Big{(}\frac{\lambda+\mu}{2}x-1\Big{)}G+\frac{(\lambda-\mu)^{2}}{4}.$
Lemma 14 and Lemma 11 imply that $(H[f])^{(d,d)}$ has co-rank at most three. Lemma 11 implies that $(H_{\lambda,\mu})^{(d,d)}$ has co-rank at most five. ∎
Proof of Theorem 5.
The theorem follows from Corollary 16 and Lemma 7. ∎
Acknowledgments. We thank Nikita Kalinin for valuable feedback on a preliminary version of this text.